# Near-Field Integrated Sensing and Communications for Secure UAV Networks

### v1

@@ III-B3 EKF-Based Data Fusion
Due to the nonlinearity of (24), we adopt the EKF for fusing the E-UAV’s measured states and the predicted ones [36]. In the $l\text{-}\mathrm{th}$ CPI, we assume that the GBS has the E-UAV’s state estimation $\hat{\boldsymbol{\mathbf{s}}}^{e}_{l-1}$ from the $(l-1)$-th CPI, along with the corresponding covariance matrix $\mathbf{C}_{l-1}$ of the estimation errors. Then, the state prediction equation and the corresponding prediction error covariance matrix are respectively given by
@@
$\mathbf{C}_{l|l-1}=\mathbf{F}\mathbf{C}_{l-1}\mathbf{F}^{T}+\mathbf{Q}_{\mathbf{s}},$ (30)
@@ IV Online Joint GBS Beamforming and C-UAV Trajectory Design Scheme
With the E-UAV trajectory tracking scheme proposed in Section III, the GBS can obtain the real-time E-UAV’s locations, and thus facilitates the online GBS beamforming and C-UAV trajectory design. On this basis, we first formulate the optimization problem with the aim of maximizing the secrecy rate for each CPI as follows:
@@ IV-A GBS Beamforming Optimization
where $[\mathbf{C}_{l}]_{m,m}$ denotes the $m$-th diagonal element of $\mathbf{C}_{l}$. To further simplify the non-convex form of (40), we utilize the Schur complement [37] to obtain the following equivalent constraint:
@@
where $\mathbf{e}_{m}\in\mathbb{R}^{6\times 1}$ denotes a vector whose $m$-th element is $1$ and other elements are 00.
@@
The relaxed problem (46), structured as a standard convex semi-definite program (SDP), can be effectively addressed by using optimization tools like CVX [39]. The specific details of the proposed algorithm for solving the original problem (35) are outlined in Algorithm 2. Note that Algorithm 2 consists of two loops. In the inner loop, the variable $\mathbf{W}_{c,l},\mathbf{W}_{e,l}$ are optimized by solving the relaxed convex form (46) iteratively. In the outer loop, the penalty factor is updated until the rank-one constraint is satisfied. The specific steps of the proposed algorithm are outlined in Algorithm 2.
@@ IV-B C-UAV Trajectory Optimization
Since the steering vector $\boldsymbol{\alpha}(r^{c}_{l},\theta^{c}_{l},\phi^{c}_{l})$ is highly complex w.r.t. $\mathbf{q}_{c,l}$, we propose to adopt the value of $\mathbf{q}_{c,l}$ in the $(k-1)$-th iteration of the SCA, i.e., $\mathbf{q}_{c,l}^{(k-1)}$, for approximating $\boldsymbol{\alpha}(r^{c}_{l},\theta^{c}_{l},\phi^{c}_{l})$ in the $k$-th iteration [40]. As such, the communication channel $\mathbf{h}^{c}_{l}$ in the $k$-th iteration of the SCA is approximated by
@@
×$\times\frac{\sigma_{c}^{2}A^{c,(k-1)}_{l}(\|\mathbf{q}_{c,l}-\mathbf{q}_{b}\|^{2}-\|\mathbf{q}_{c,l}^{(k)}-\mathbf{q}_{b}\|^{2})}{(B^{c,(k-1)}_{l,}+\gamma^{(k)})^{2}+A^{c,(k-1)}_{l}(B^{c,(k-1)}_{l}+\gamma^{(k)})}$
@@ IV-C Overall Algorithm
$N$ Number of symbols in one CPI $500$ $\mathbf{q}_{\mathrm{min}}$ Flight area lower bound $\left[-30,-30,25\right]^{T}$
@@
$\Delta t$ CPI duration $0.05$ s $V_{\mathrm{max}}$ Maximum velocity of C-UAV $5$ m/s
$\beta_{0}$ Channel power gain at $1$ meter $-30$ dB $d_{\mathrm{min}}$ Minimum safe distance $7$ m
$P_{\text{max}}$ GBS maximum transmit power $30$ dBm $\sigma_{c}^{2}$,$\sigma_{e}^{2}$ C-UAV and E-UAV noise power $-80$ dBm
$\sigma_{\mathrm{RCS}}$ Radar cross-section of the E-UAV $0.03$ $\mathrm{m}^{2}$ $\sigma_{b}^{2}$ GBS noise power $-50$ dBm
$c_{r}$,$c_{\theta}$,$c_{\phi}$ Scaling factors for location measurement errors $10^{-6}$,$10^{-2}$,$10^{-2}$ $c_{v_{r}},c_{v_{\theta}},c_{v_{\phi}}$ Scaling factors for velocity measurement errors $1$,$1$,$1$
@@ V-B E-UAV Trajectory Tracking and C-UAV Trajectory Design Performance
In Fig. 6, we depict the estimated E-UAV trajectory obtained with the trajectory tracking scheme designed in Section III as well as the C-UAV trajectory obtained with Algorithm 4 given different initial locations. Specifically, we set three different C-UAV’s initial locations, given by $[10,2,30]^{T}$, $[14,6,30]^{T}$, $[12,16,30]^{T}$, and label the corresponding trajectory as “C-UAV trajectory 1”, “C-UAV trajectory 2”, and “C-UAV trajectory 3”, respectively. The E-UAV’s ground-truth trajectory is also plotted for comparison. One can first observe that estimated trajectory perfectly aligns with the ground-truth trajectory of the E-UAV, which underscores the high accuracy for tracking the E-UAV trajectory with the proposed near-field 3D sensing method. Observing the C-UAV’s trajectories, one can find that the C-UAV’s flight tends to approach the GBS, which is expected as the secrecy rate can be improved when the path loss between the GBS and C-UAV is reduced. Moreover, it is also seen that the C-UAV keeps adjusting its flight during the given time period rather than adopting the straight-line trajectory. The is because, due to the presence of the E-UAV, the C-UAV trajectory should be designed to guarantee the minimum safe distance, as well as to reduce the channel correlations with the E-UAV so as to improve the secrecy rate. From the 3D trajectory plot given in Fig. 6(b), we also see that the C-UAV initially descends rapidly along the z-axis until reaches to the predetermined minimum flight altitude, after which the height is not changed anymore. This is quite intuitive as the C-UAV tends to be get close to the GBS to improve the achievable rate.
@@
In Fig. 7, we plot the instantaneous normalized E-UAV’s state estimation error (NSEE) during the whole flight duration of $80$ CPIs. The NSEE is defined as $\mathrm{log_{10}(\frac{|\mathbf{s}^{e}_{l}-\hat{\mathbf{s}}^{e}_{l}|^{2}}{\rho})}$, with the normalization factor $\rho$ set to $10$. It is evident to see that the NSEE increases with the increment of the maximum E-UAV tracking MSE $\Gamma$. Specifically, define the average NSEE over $L$ CPIs as $\frac{1}{L}\sum_{l=1}^{L}\mathrm{log_{10}(\frac{|\mathbf{s}^{e}_{l}-\hat{\mathbf{s}}^{e}_{l}|^{2}}{\rho})}$. We obtain the average NSEE over the $80$ CPIs as $-1.09$ dB, $-1.98$ dB, $-2.18$ dB, and $-2.58$ dB for $\Gamma$ given by $0.2$, $0.05$, $0.04$, and $0.02$, respectively. This is expected, as the instantaneous E-UAV’s state estimation error gets smaller when the estimation accuracy requirement becomes more strict. We also observe that the instantaneous NSEE provided in Fig. 7 does not monotonically decrease over time, which is due to the E-UAV’s varying locations during the whole time period.

### v2

@@ III-B3 EKF-Based Data Fusion
Due to the nonlinearity of (24), we adopt the EKF for fusing the E-UAV’s measured states and the predicted ones [39]. In the $l\text{-}\mathrm{th}$ CPI, we assume that the GBS has the E-UAV’s state estimation $\hat{\boldsymbol{\mathbf{s}}}^{e}_{l-1}$ from the (l1)(l-1)-th CPI, along with the corresponding covariance matrix $\mathbf{C}_{l-1}$ of the estimation errors. Then, the state prediction equation and the corresponding prediction error covariance matrix are respectively given by
@@
$\mathbf{C}_{l|l-1}=\mathbf{F}\mathbf{C}_{l-1}\mathbf{F}^{T}+\mathbf{Q}_{s},$ (30)
@@ IV Online Joint GBS Beamforming and C-UAV Trajectory Design Scheme
With the E-UAV trajectory tracking scheme proposed in Section III, the GBS can obtain the real-time E-UAV’s locations, and thus facilitates the online GBS beamforming and C-UAV trajectory design. On this basis, we formulate the optimization problem with the aim of maximizing the secrecy rate for the $l$-th CPI as follows:
@@ IV-A GBS Beamforming Optimization
where $\mathbf{C}_{l}]_{m,m}$ denotes the $m$-th diagonal element of $\mathbf{C}_{l}$. To further simplify the non-convex form of (40), we utilize the Schur complement [40] to obtain the following equivalent constraint:
@@
where $\mathbf{e}_{m}\in\mathbb{R}^{6\times 1}$ denotes a vector whose $m$-th element is $1$ and other elements are 0.
@@
The relaxed problem (46), structured as a standard convex semi-definite program (SDP), can be effectively addressed by using optimization tools like CVX [42]. The specific details of the proposed algorithm for solving the original problem (35) are outlined in Algorithm 2. Note that Algorithm 2 consists of two loops. In the inner loop, the variable $\mathbf{W}_{c,l},\mathbf{W}_{e,l}$ are optimized by solving the relaxed convex form (46) iteratively. In the outer loop, the penalty factor $\omega$ is updated with the scaling factor $c>1$ until the rank-one constraint is satisfied.
@@ IV-B C-UAV Trajectory Optimization
Since the steering vector $\boldsymbol{\alpha}(r^{c}_{l},\theta^{c}_{l},\phi^{c}_{l})$ is highly complex w.r.t. $\mathbf{q}_{c,l}$, we propose to adopt the value of $\mathbf{q}_{c,l}$ in the (k1)(k-1)-th iteration of the SCA, i.e., $\mathbf{q}_{c,l}^{(k-1)}$, for approximating $\boldsymbol{\alpha}(r^{c}_{l},\theta^{c}_{l},\phi^{c}_{l})$ in the $k$-th iteration [43]. As such, the communication channel $\mathbf{h}^{c}_{l}$ in the $k$-th iteration of the SCA is approximated by
@@
$\times\frac{\sigma_{c}^{2}A^{c,(k-1)}_{l}(\|\mathbf{q}_{c,l}-\mathbf{q}_{b}\|^{2}-\|\mathbf{q}_{c,l}^{(k)}-\mathbf{q}_{b}\|^{2})}{(B^{c,(k-1)}_{l,}+\gamma^{(k)})^{2}+A^{c,(k-1)}_{l}(B^{c,(k-1)}_{l}+\gamma^{(k)})}$
@@ IV-C Overall Algorithm
NN Number of symbols in one CPI $500$ $\mathbf{q}_{\mathrm{min}}$ Flight area lower bound $\left[-30,-30,25\right]^{T}$
@@
$\Delta t$ CPI duration $0.05$ s $V_{\mathrm{max}}$ Maximum velocity of C-UAV 55 m/s
$\beta_{0}$ Channel power gain at $1$ meter 30-30 dB $d_{\mathrm{min}}$ Minimum safe distance 77 m
$P_{\text{max}}$ GBS maximum transmit power $30$ dBm $\sigma_{c}^{2}$$\sigma_{e}^{2}$ C-UAV and E-UAV noise power 80-80 dBm
$\sigma_{\mathrm{RCS}}$ Radar cross-section of the E-UAV $0.03$ $\mathrm{m}^{2}$ $\sigma_{b}^{2}$ GBS noise power 50-50 dBm
[…]
@@ V-B E-UAV Trajectory Tracking and C-UAV Trajectory Design Performance
In Fig. 6, we depict the estimated E-UAV trajectory obtained with the trajectory tracking scheme designed in Section III as well as the C-UAV trajectory obtained with Algorithm 4 given different initial locations. Specifically, we set three different C-UAV’s initial locations, given by $10,2,30]^{T}$, $14,6,30]^{T}$, $12,16,30]^{T}$, and label the corresponding trajectory as “C-UAV trajectory 1”, “C-UAV trajectory 2”, and “C-UAV trajectory 3”, respectively. The E-UAV’s ground-truth trajectory is also plotted for comparison. One can first observe that estimated trajectory perfectly aligns with the ground-truth trajectory of the E-UAV, which underscores the high accuracy for tracking the E-UAV trajectory with the proposed near-field 3D sensing method. Observing the C-UAV’s trajectories, one can find that the C-UAV’s flight tends to approach the GBS, which is expected as the secrecy rate can be improved when the path loss between the GBS and C-UAV is reduced. Moreover, it is also seen that the C-UAV keeps adjusting its flight during the given time period rather than adopting the straight-line trajectory. The is because, due to the presence of the E-UAV, the C-UAV trajectory should be designed to guarantee the minimum safe distance, as well as to reduce the channel correlations with the E-UAV so as to improve the secrecy rate. From the 3D trajectory plot given in Fig. 6(b), we also see that the C-UAV initially descends rapidly along the z-axis until reaches to the predetermined minimum flight altitude, after which the height is not changed anymore. This is quite intuitive as the C-UAV tends to be get close to the GBS to improve the achievable rate.
@@
In Fig. 7, we plot the instantaneous normalized E-UAV’s state estimation error (NSEE) during the whole flight duration of $80$ CPIs. The NSEE is defined as $\mathrm{log_{10}(\frac{|\mathbf{s}^{e}_{l}-\hat{\mathbf{s}}^{e}_{l}|^{2}}{\rho})}$, with the normalization factor $\rho$ set to $10$. It is evident to see that the NSEE increases with the increment of the maximum E-UAV tracking MSE $\Gamma$. Specifically, define the average NSEE over $L$ CPIs as $\frac{1}{L}\sum_{l=1}^{L}\mathrm{log_{10}(\frac{|\mathbf{s}^{e}_{l}-\hat{\mathbf{s}}^{e}_{l}|^{2}}{\rho})}$. We obtain the average NSEE over the $80$ CPIs as 1.09-1.09 dB, 1.98-1.98 dB, 2.18-2.18 dB, and 2.58-2.58 dB for $\Gamma$ given by $0.2$, $0.05$, $0.04$, and $0.02$, respectively. This is expected, as the instantaneous E-UAV’s state estimation error gets smaller when the estimation accuracy requirement becomes more strict. We also observe that the instantaneous NSEE provided in Fig. 7 does not monotonically decrease over time, which is due to the E-UAV’s varying locations during the whole time period.